OpenStudy (anonymous):

convert y''+2ky'+a^2 y=0 into 2*2 system please?

OpenStudy (michele_laino):

here we can introduce a new function \(\Large \eta \), so, we can rewrite your system as below: \[\huge \left\{ \begin{gathered} y' = \eta \hfill \\ \eta ' + 2k\eta + {a^2}y = 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

i need it as a 2*2matrix form

OpenStudy (irishboy123):

you have it now

OpenStudy (anonymous):

i don't understand it

OpenStudy (anonymous):

can you write it as a matrix please

OpenStudy (irishboy123):

let's have a go!

OpenStudy (anonymous):